OpenStudy (anonymous):

Input the equation of the given line in standard form. The line including (3, 1) and (-2, 3).

OpenStudy (anonymous):

y-y1=m(x-x1) m=(y2-y1)/(x2-x1) gave you the formulas now solve them or do you need help

OpenStudy (anonymous):

(2,-5)?